OpenStudy (anonymous):

log2=a and log3=b. what is x in 3^(2^x)=2^(3^x)

geerky42 (geerky42):

take log of both side \[\Large 3^{(2^x)} = 2^{(3^x)} \\\Large \ln3^{(2^x)} = \ln 2^{(3^x)} \\\Large 2^x\ln3 = 3^x\ln 2\] Now divide both sides by \(3^x\ln3\) to get \(\large\dfrac{2^x}{3^x} =\left(\dfrac{2}{3}\right)^x = \dfrac{\ln2}{\ln3}\) Now take log again, you should be able to solve for x from here